Use Poisson Distribution to solve problems 6-7 6. Suppose that the average number of accidents occurring weekly on a particular stretch of a highway equals 2. What is the probability that within next week: a) 0 accidents occur P(x 0) (3 points) A) 0.1258 B) 0.1353 C) 0.8647 D) 0.2706 b) 1 or less accidents occur P( (5 points) 2)-
Answer #1

Solution:
Given in the question lambda = 2
We need to find P(X=0)
Probability of Poission distribution can be calculated as
P(X=0) = e^(-lambda) * lambda^x / X! = e^-2 * (2^0)/0! = 0.1353
So its answer is B.i.e. 0.1353

Solution(b)
P(X<2) = P(X=0) + P(X=1) = 0.1353 + e^(-1) * (2^1)/1! = 0.1353 + (0.3678*2) = 0.1353 + 0.7356 = 0.8709